At the half-equivalence point, the pH = pKa. Based on your average above, what is the value of Ka for the acid? Ka = ___ M

PH= 4.86 for the avg
Answer #1

Given that

pH = 4.86

So, pKa = pH = 4.86

Ka = 10–pKa = 10–4.86 = 1.38038 × 10–5 = 1.38 × 10–5

Ka for the acid = 1.38 × 10–5